OpenStudy (anonymous):

if p arithmetic means are inserted between a nad b ,prove that d=b-a/p+1

OpenStudy (amistre64):

mightbe something akin to this let pn = {p1,p2,p3,p4,...,pn} p1 = (a+p2)/2 p2 = (p1+p3)/2 = (a+p2+p3)/2 p3 = (p2 + p4)/2 = (a+p2+p3+p4)/2 ... pn = (a+p2+p3+p4+...+pn+b)/2 2pn = a+p2+p3+p4+...+pn+b 2pn - pn = a+p2+p3+p4+...+p{n-1}+b pn(2 - 1) = a+p2+p3+p4+...+p{n-1}+b pn = a+p2+p3+p4+...+p{n-1}+b

OpenStudy (amistre64):

but then maybe not :/